Solve the system by addition.
2x + 4y = 3
4x – 2y = 1

A) (1/2, 1/2) B) (1/3, 1/3) C) (1/4, 1/4) D)(1/5, 1/5)

Answer by scott8148(6628)

multiplying the second equation by 2 and adding the result to the first equation gives (2x+4y)+2(4x-2y)=3+2(1)

2x+4y+8x-4y=3+2 ... 10x=5 ... x=1/2 ... 2(1/2)+4y=3 ... 1+4y=3 ... 4y=2 ... y=1/2 ... answer A)